The tip of an airplane propeller blade is travelling at 295 m/s. If the radius of the propeller is 1.35 m, what is the period of rotation of the propeller blade?
1
Expert's answer
2020-07-27T09:34:46-0400

The speed of an end of airplane propeller blade is related with its period of rotation by equation

"v=\frac{2\pi R}{T}"

Hence, the period of rotation of the propeller blade


"T=\frac{2\pi R}{v}=\frac{2\pi \times 1.35\:\rm m}{295\:\rm m/s}=0.0287\:\rm s"